### v1

@@ Abstract
The global ocean model NEMO is run in a series of stand-alone configurations (2015–2022) to investigate the potential for improving global medium-range storm surge forecasts by including the inverse barometer effect. Here, we compare a control experiment, where the inverse barometer effect was not included, against a run dynamically forced with mean sea level pressure. In the control experiment, the inverse barometer effect was then calculated diagnostically and added to the ocean model sea surface elevation, resulting in a total of three experiments to investigate. We compare against the global GESLA3 water level data set and find that the inclusion of the inverse barometer effect reduces the root-mean-square error by $\sim 1$ cm on average. When we mask out all data where the observed storm surge is less than $\pm 1$ or $\pm 2$ standard deviations, including the inverse barometer effect reduces the RMS error by 454-5 cm. While both methods reduce water level errors, there are regional differences in their performance. The run with dynamical pressure forcing is seen to perform slightly better than diagnostically adding the inverse barometer effect in enclosed basins such as the Baltic Sea. Finally, an ensemble forecast experiment with the Integrated Forecast System of the European Centre for Medium-range Weather Forecasts demonstrates that when the diagnostic inverse barometer effect is included for a severe storm surge event in the North Sea (Storm Xaver, December 2013), the ensemble spread of water level provides a stronger and earlier indication of the observed maximum surge level than the when the effect is excluded.
@@ 1 Introduction
The mean sea level varies on time scales ranging from hours (tides and synoptic weather) to decades (climate change and large-scale climate fluctuations). In addition to the steric changes (expansion and contraction due to changes in the heat content and the salinity of the ocean interior), changes to the meltwater input affect the long-term trends in still water level. On much shorter time scales, we have the synoptic fluctuations caused by variations in the sea level pressure of the overlying atmosphere, known as the inverse barometer effect (IB). This effect, where a decrease in the surface pressure of 1 hPa leads to a local increase in water level of approximately 1 cm is a practical rule of thumb allowing quick estimates of the impact of atmospheric pressure on the water level (Piecuch and Ponte, 2015). In addition, frictional effects will lead to water level variations that have a much wider geographical extent. The simplest model on a rotating sphere is the Ekman transport (Gill, 1982)[pp 319–322], where a balance is reached between the frictional force of the wind on the sea surface and the Coriolis effect. When the water is pushed against a coastline, or drawn away from the coastline, the effect typically sets up Kelvin waves trapped by the coast. Such water level variations can be substantial, and are typically among the most devastating effects of tropical cyclones. The extratropics also experience large fluctuations due to Ekman effects, dramatically demonstrated by, among other cases, the devastating storm surge of 1953 in the North Sea (Wolf and Flather, 2005). Storm surge modeling is a routine activity of national hydro-meteorological services (Kristensen et al., 2022) and can be carried out with modeling systems of varying degrees of complexity, ranging from two-dimensional homogeneous equations of the vertically averaged (barotropic) current and its associated water level to more complex fully baroclinic three dimensional ocean models. Such models are run on regional scales and in some cases globally. The models can include the major constituents of the astronomical tide, or these can be added diagnostically. Regional forecast systems typically rely on boundary conditions from basin-scale or global model systems. Their horizontal resolution is high enough (4 km is a typical number) that local topographic effects can be accounted for. Since water level is mostly a concern along coastlines (but not just, fixed platforms must account for changes in the water level, and floating structures may have to account for tension on moorings), such forecast systems are expected to yield reasonably accurate water level predictions in semi-enclosed waters. Furthermore, the uncertainty associated with the passage of synoptic weather systems translates into uncertainty of the water level they cause. Ensemble prediction is thus a natural approach for water level forecasting.
@@
Here we present a comparison of two standalone global NEMO (Madec et al., 2024) ocean model integrations with and without the atmospheric pressure effect, forced with the atmospheric reanalysis ERA5 (Hersbach et al., 2020). No astronomical tidal constituents are included in the model setup. The model runs cover the period 2015–2022, inclusively. The purpose of our study is two-fold. First, we want to investigate how well a coarse (roughly $0.25^{\circ}$) fully baroclinic ocean model can be expected to represent coastal and open-ocean water level variations associated with the passage of tropical and extratropical cyclones. Secondly, we want to assess how well a model without the inverse barometer effect (only Ekman effects) can be expected to perform. This we do by diagnostically adding the inverse barometer effect to a control experiment with constant sea level pressure. The model is identical to the ocean model component of the Integrated Forecast System (IFS) of the European Centre for Medium-range Weather Forecasts (ECMWF). See Mogensen et al. 2012 and Mogensen et al. 2017 for further details on the ocean model component of IFS.
@@
Two experiments have been performed to assess two alternative ways of including the IB effect and its impact on the modeled sea surface elevation. This leaves us with three variables representing the water level, hereafter referred to as noIB, the control experiment with no IB effect, diaIB, diagnostically adding the IB effect to the control experiment, and finally dynIB, the water level from NEMO including the dynamically computed effect of the sea level pressure. Wind stress, and its effect on the water level, is included in all three variables.
@@
The investigation illustrates the potential for local storm surge modelling, by exploring the impact in the global IFS ensemble, which is commonly used to provide boundary conditions for high resolution Limited Area Models (LAMs). It also highlights the usefulness of a global, relatively coarse, coupled atmosphere-ocean model for probabilistic forecasting and early warning of storm surge along complex coastlines.
@@ 2 Observational dataset
The observation dataset used in this work is the Global Extreme Sea Level Analysis version 3 (GESLA3), see Haigh et al. 2023. GESLA3 is a comprehensive and freely available dataset containing observation records from 36 different sources covering large parts of the globe. The dataset consists of 5119 stations with time series of observed total water level (TWL) on a common data format, covering the period from 1800 until 2020. It contains an aggregated total of 91,021 years for all the records. We note that some observation stations are contained in more than one of the sources used to compile the dataset, hence the total number of unique observation stations is less than the number of records. The locations and geographical distribution of the records are shown in Figure 1. To isolate the weather-induced water level (hereafter referred to as the storm surge for brevity) from the TWL, we performed a tidal analysis of each record using the 12 dominant tidal constituents in the pangeo-pytide python package, based on the FES2014 global ocean tide atlas (Lyard et al., 2021), and subtracted this tidal signal from the observed TWL. The residual signal is hereafter referred to as the observed storm surge. Before the analysis, the temporal mean for both model time series and observation records is removed. We utilize all the records that overlap in time with our simulation period, located no more than two grid points away from an ocean grid point in the model. This reduces the number of records to be used in the analysis to around 1000 records. The blue dots in Figure 1 indicate the locations and geographical distribution of the records used in our analysis, while the red dots indicate the records that was not used. All observation records have been resampled to hourly time resolution through linear interpolation to facilitate the comparison with ocean model data. This was done after the tidal analysis was performed.
@@
A feature of the GESLA3 dataset is that some stations appear as multiple records in the dataset, i.e., are provided from multiple sources, like, e.g., the national hydro-meteorological services and the Pan-European Copernicus Marine Service (CMEMS). These seem to occasionally be resampled or interpolated from their original temporal resolution to a different resolution, often hourly. When performing the tidal analysis on these records, and subtracting the tidal signal from the TWL, we are often left with a very noisy storm surge signal. This tends to clutter the model evaluation presented in Section 4, and even if our brief investigation into the issue suggests that the number of “noisy” data points is small, this feature could have an impact on the general statistics when such records are included. More details and a visual example can be found in Appendix A: Errors in observations. Attempts to remove these outliers has not been successful, and we are therefore left with the choice of either omitting the records entirely, or leaving them in the dataset. We have chosen the latter.
@@ 3 Model setup
The NEMO ocean model (Madec et al., 2024) is here run with atmospheric forcing from the ERA5 reanalysis (Hersbach et al., 2020). The domain is a global stretched tripolar ORCA025 grid with an average resolution of $0.25\times 0.25^{\circ}$ which is refined threefold in the Equatorial band. The vertical is resolved with a $z$ coordinate with variable resolution. The upper level is $1\,\mathrm{m}$ thick. This setup is commonly referred to as ORCA025_Z75. The ocean model reads the 10 m wind vectors, temperature, humidity, heat and water fluxes from ERA5 every hour. The ocean model also reads forcing from the ECWAM wave model (a component of ERA5) as described by Breivik et al. 2015. We now perform an ocean-only integration (experiment id i5k6) covering the period 2015–2022 (from a warm start). We then compute the inverse barometer effect from the ERA5 mean sea level pressure (MSLP) and output the total sea surface elevation (including the inverse barometer effect) as well as the prognostic elevation (only frictional and baroclinic effects). This control run is then compared against an otherwise identical ocean-only run where the ocean model is allowed to respond to a varying MSLP taken from ERA5 (experiment id i5k8). This run is referred to as the dynamical experiment (dynIB) whereas the former is referred to as the diagnostic experiment (diaIB) (since we compute the inverse barometer effect diagnostically from the local MSLP). Neither run includes tides. The variables of interest in the control run are thus the sea surface elevation without the inverse barometer effect, referred to as noIB, and the sea surface elevation including the diagnostically computed inverse barometer effect, diaIB. These are compared against the total sea surface elevation dynIB in the alternative model setup with dynamical MSLP.
In Section 6, we use the IFS ensemble prediction system (ENS, e.g. ECMWF, 2024a) to showcase how inclusion of the IB effect can impact the probabilistic forecast of storm surges. ENS is a coupled global forecasting system, which uses NEMO for its ocean model component, thus enabling us to explore how the impacts seen from the standalone NEMO studies translate into a coupled forecasting scenario.
Both the uncoupled and coupled forecasts use NEMO version 4.0, which is a newer version than the 3.4 version of NEMO which is currently in operational use at ECMWF. The coupled model runs were based on a development version of IFS CY49R2 to incorporate the changes needed to run the coupled setup with NEMO V4.0.
@@ 4.2 Model evaluation and validation
The control experiment with diagnostic inclusion of the IB effect (diaIB) and the alternative (dynIB) is compared with the GESLA3 observations. As a baseline, we also compare against the control experiment without inclusion of the IB effect (noIB). We focus the evaluation on the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE), together with scaled versions of the two ($\mathrm{MAE}/\sigma$ and $\mathrm{RMSE}/\sigma$, where $\sigma$ is the standard deviation of the observed storm surge). The values of MAE and RMSE for each record is scaled by the standard deviation of the observed storm surge for that record to provide a measure of the relative error, since the magnitude of the absolute error tends to be correlated with the magnitude of the storm surge, to allow us to compare the model performance across different records with different typical amplitudes. This is in line with the approach by Kristensen et al. 2024, and the formulas can be found in Appendix B.
The MAE and RMSE [m] for the water level of noIB, diaIB and dynIB are listed in Table 1. We focus on the median of the error metrics across the records, but also include the mean values. The median is chosen as it is a robust estimate of central tendency and thus suitable for analysis given the large outliers found in some of the observation records, as explained in Section 2. We include the MAE and RMSE for the entire hindcast period, denoted by the subscript all, and also values where we mask the data when the observed absolute values of the storm surge is less than one and two standard deviations from the mean, denoted by the subscripts $\pm 1\mathrm{std}$ and $\pm 2\mathrm{std}$. This is done to emphasize the cases and time periods when the contribution from the storm surge to the total water level is significant to focus on the model’s ability to capture anomalously low and high water level events (the latter being what we usually think of as storm surges). Similarly, the relative (scaled) values $\mathrm{MAE}/\sigma$ and $\mathrm{RMSE}/\sigma$, are presented in Table 2.
@@
However, Figure 3 shows that the statistical distribution of the RMSE has a large variation, with the majority of the errors in the range $4-30\,\mathrm{cm}$ and $0.3-1.6]\sigma$, when considering all data. It is evident that all the distributions, except the relative noIB, are skewed to the right. This is as expected based on the mean and median values in Table 1 and 2 where the mean value is generally larger than the median. The width of the distributions increases with stricter masking criteria.
@@
Figure 3: Distribution of the Root Mean Square Error (RMSE) for all records for the three different storm surge representations. The top panels (a-c) present the RMSE in meters, with a bin size of 2 cm. The bottom panels (d-f) show the RMSE scaled by the standard deviation ($\sigma$) of each observation record, using a bin size of $0.1\sigma$. Panels (a) and (d) present the RMSE distribution for the complete dataset. Panels (b) and (e) illustrate the RMSE for observed storm surge absolute values more than one standard deviation ($|\mathrm{obs}|>\sigma$). Panels (c) and (f) show the RMSE for observed storm surge absolute values exceeding two standard deviations ($|\mathrm{obs}|>2\sigma$). The $y$-axis indicates the number of occurrences (records).
Additionally, the map in Figure 4 clearly shows the regional differences in performance between the two experiments. The difference in relative error, displayed in Figure 4, highlights the European NWS and the Baltic Sea as two regions (see also Figure 5) where the differences between the two experiments are large with diaIB outperforming dynIB on the European NWS, and vice versa in the Baltic Sea. Most other areas exhibit relatively small differences, with a few scattered exceptions mostly in favor of the diaIB. It is unclear why the dynIB experiment seems to outperform the diagnostic experiment in the semi-enclosed water body of the Baltic Sea. However, as we see a similar, albeit weaker, tendency in the Mediterranean Sea and the Black Sea, we speculate that the narrow connections to the open ocean limits the efficacy of the inverse barometer effect. These water bodies also have very small tidal amplitudes for the same reason, see e.g., Medvedev et al. 2016. In other words, although the atmospheric pressure variations set up a pressure gradient, the narrows connecting the Baltic Sea (Weisse et al., 2021), the Mediterranean Sea and the Black Sea (respectively the Danish Straits, the Strait of Gibraltar and the Bosphorus) are not wide enough for water to adjust on the synoptic time scales of hours to at most a few days. When the atmospheric pressure averaged over the entire confined sea area changes, this will in the diaIB lead to a change in the average water level, whereas in reality and in the dynIB, these changes in water level are mostly local, leaving the water level over the confined sea area mostly unchanged during the passage of extratropical systems. This effect becomes more pronounced when the dimensions of the confined water body coincides with the spatial extent of the weather systems.

### v2

@@ Abstract
The global ocean model NEMO is run in a series of stand-alone configurations (2015–2022) to investigate the potential for improving global medium-range storm surge forecasts by including the inverse barometer effect. The analysis focus on the residual water level, i.e. the water level variations not due to tides. Here, we compare a control experiment, where the inverse barometer effect was not included, against a run dynamically forced with mean sea level pressure. In the control experiment, the inverse barometer effect was then calculated diagnostically and added to the ocean model sea surface elevation, resulting in a total of three experiments to investigate. We compare against the global GESLA3 water level data set and find that the inclusion of the inverse barometer effect reduces the root-mean-square error by $\sim 1$ cm on average. When we mask out all data where the observed residual water level is less than $\pm 1$ or $\pm 2$ standard deviations, including the inverse barometer effect reduces the RMS error by 454-5 cm. While both methods reduce water level errors, there are regional differences in their performance. The run with dynamical pressure forcing is seen to perform slightly better than diagnostically adding the inverse barometer effect in enclosed basins such as the Baltic Sea. Finally, an ensemble forecast experiment with the Integrated Forecast System of the European Centre for Medium-range Weather Forecasts demonstrates that when the diagnostic inverse barometer effect is included for a severe storm surge event in the North Sea (Storm Xaver, December 2013), the ensemble spread of water level provides a stronger and earlier indication of the observed maximum surge level than the when the effect is excluded.
@@ 1 Introduction
The mean sea level varies on time scales ranging from hours (tides and synoptic weather) to decades (climate change and large-scale climate fluctuations) (gregory2019). In addition to the steric changes (expansion and contraction due to changes in the heat content and the salinity of the ocean interior), changes to the meltwater input affect the long-term trends in still water level. On much shorter time scales, we have the synoptic fluctuations caused by variations in the sea level pressure of the overlying atmosphere, known as the inverse barometer (IB) effect (wunsch1997). While typically isostatic, dynamic behavior arises at specific wave frequencies. ponte1991 found that global departures from the pure IB response are usually small ($1-3~\mathrm{cm}$), but increase to $5\%-20\%$ for periods between 2 and 7 days. Furthermore, pressure at high latitudes can drive sea level changes at lower latitudes, a nonlocal effect that local IB (diagnostic) corrections miss. These studies demonstrate that dynamic forcing is essential to accurately model high-frequency water level changes.
For the IB effect, a decrease in the surface pressure of 1 hPa leads to a local increase in water level of approximately 1 cm is a practical rule of thumb allowing quick estimates of the impact of atmospheric pressure on the water level (piecuch2015). In addition, frictional effects will lead to water level variations that have a much wider geographical extent. The simplest model on a rotating sphere is the Ekman transport (gil82)[pp 319–322], where a balance is reached between the frictional force of the wind on the sea surface and the Coriolis effect. When the water is pushed against a coastline, or drawn away from the coastline, the effect typically sets up Kelvin waves trapped by the coast. Such water level variations can be substantial, and are typically among the most devastating effects of tropical cyclones. The extratropics also experience large fluctuations due to Ekman effects, dramatically demonstrated by, among other cases, the devastating storm surge of 1953 in the North Sea (wolf05). We note that Ekman transport and the IB effect does not always act simply as additive components, but can sometimes be dynamically coupled, as shown by gomis2008. They show how, in certain regions, wind forcing can sometimes counteract and override the IB response.
@@
Storm surge modeling is a routine activity of national hydro-meteorological services (see e.g. kristensen22) and can be carried out using modeling systems of varying degrees of complexity, ranging from two-dimensional homogeneous equations of the vertically averaged (barotropic) current and its associated water level to more complex fully baroclinic three dimensional ocean models. Such models are run on regional scales and in some cases globally (cirano2025). The models can include the major constituents of the astronomical tide, or these can be added diagnostically. Regional forecast systems typically rely on boundary conditions from basin-scale or global model systems. Their horizontal resolution is high enough (around 4 km is a typical number) that local topographic effects can be accounted for. Since variations in water level are mostly a concern along coastlines (but not only, fixed platforms must account for changes in the water level, and floating structures may have to account for tension on moorings), such forecast systems are expected to yield reasonably accurate water level predictions in semi-enclosed waters. Furthermore, the uncertainty associated with the passage of synoptic weather systems translates into uncertainty of the water level they cause. Ensemble prediction is thus a natural approach for water level forecasting.
In addition to dedicated storm surge models, any ocean model with a free-surface, which includes most global and regional ocean models, is capable of predicting water levels. However, most global, and some of the larger regional, models do not include the water level contribution from the IB effect (static IB assumption). The present work seeks to explore whether this shortcoming can be overcome by adding the IB effect as a post process, or if it must be included at run-time.
The limitations of the static IB assumption, specifically its failure at periods shorter than 3 days and in high-latitude regions, have been well-documented. carrere2003 demonstrated that a global barotropic model (MOG2D-G) using a finite element unstructured mesh (ranging from 400 km in the deep ocean to 20 km in coastal areas) could reduce sea level variance by more than 50% at tide gauge locations compared to the classical IB correction. A critical finding of their work was that while dynamic pressure forcing alone offered modest improvements, the inclusion of wind stress was responsible for the majority of the variance reduction.
We illustrate the potential for local storm surge modelling, by exploring the impact in the global Integrated Forecast System (IFS) ensemble of the European Centre for Medium-range Weather Forecasts (ECMWF), which is commonly used to provide boundary conditions for high resolution Limited Area Models (LAMs). We also highlight the usefulness of a global, relatively coarse, coupled atmosphere-ocean model for probabilistic forecasting and early warning of storm surge along complex coastlines.
Here we present a comparison of two standalone global NEMO (gurvan_madec_2024) ocean model integrations covering the period 2015–2022, inclusively, with and without the atmospheric pressure effect, forced with the atmospheric reanalysis ERA5 (hersbach20era5). The model is identical to the ocean model component of the ECMWF IFS. See mogensen12 and mogensen17 for further details on the ocean model component of IFS. No astronomical tidal constituents are included in the model setup. This means that the water level given by the model is only the atmospheric contribution to the total water level. It is comparable to the residual water level due to atmospheric effects in the observations, i.e. the residual when subtracting the astronomical tides from the observed total water level, not taking non-linear interactions between tides and the atmospheric effects into account.
@@
The purpose of our study is two-fold. First, we want to investigate how well a coarse (roughly $0.25^{\circ}$) fully baroclinic ocean model can be expected to represent coastal and open-ocean water level variations associated with the passage of tropical and extratropical cyclones. Secondly, we want to assess how well a model without the inverse barometer effect (only Ekman effects) can be expected to perform.
@@
Two experiments have been performed to assess two alternative ways of including the IB effect and its impact on the modeled sea surface elevation. This leaves us with three variables representing the water level, referred to as noIB, the control experiment with no IB effect, diaIB, diagnostically adding the IB effect to the control experiment, and finally dynIB, the water level from NEMO including the dynamically computed effect of the sea level pressure. Wind stress, and its effect on the water level, is included in all three variables.
@@ 2 Observational dataset
The observation dataset used in this work is the Global Extreme Sea Level Analysis version 3 (GESLA3), see haigh2023. GESLA3 is a comprehensive and freely available dataset containing observation records from 36 different sources covering large parts of the globe. The dataset consists of 5119 stations with time series of observed total water level (TWL) on a common data format, covering the period from 1800 until 2020. It contains an aggregated total of 91,021 years for all the records. We note that some observation stations are contained in more than one of the sources used to compile the dataset, hence the total number of unique observation stations is less than the number of records. The locations and geographical distribution of the records are shown in Figure 1. To isolate the weather-induced effects on the water level, i.e. the residual water level, from the TWL, we performed a tidal analysis of each record using 11 of the dominant tidal constituents [^1] The tidal constituents used in the de-tiding are: M2, S2, N2, K1, M4, O1, MN4, Q1, P1, K2, MS4 in the pangeo-pytide python package, based on the FES2014 global ocean tide atlas (lyard21), and subtracted this tidal signal from the observed TWL. This is hereafter referred to as the observed residual water level or residual. Before the analysis, the temporal mean for both model time series and observation records is removed. We utilize all the records that overlap in time with our simulation period, located no more than two grid points away from an ocean grid point in the model. This reduces the number of records to be used in the analysis to around 1000 records. The blue dots in Figure 1 indicate the locations and geographical distribution of the records used in our analysis, while the red dots indicate the records that was not used. All observation records have been resampled to hourly time resolution through linear interpolation to facilitate the comparison with ocean model data. This was done after the tidal analysis was performed.
@@
A feature of the GESLA3 dataset is that some stations appear as multiple records in the dataset, i.e., are provided from multiple sources, like, e.g., the national hydro-meteorological services and the Pan-European Copernicus Marine Service (CMEMS). These seem to occasionally be resampled or interpolated from their original temporal resolution to a different resolution, often hourly. When performing the tidal analysis on these records, and subtracting the tidal signal from the TWL, we are often left with a very noisy residual signal. This tends to clutter the model evaluation presented in Section 4, and even if our brief investigation into the issue suggests that the number of “noisy” data points is small, this feature could have an impact on the general statistics when such records are included. More details and a visual example can be found in Appendix A. Attempts to remove these outliers have not been successful, and we are therefore left with the choice of either omitting the records entirely, or leaving them in the dataset. We have chosen the latter.
@@ 3 Model setup
3.1 Model description
The NEMO ocean model (gurvan_madec_2024) is here run with hourly atmospheric forcing from the ERA5 reanalysis (hersbach20era5). The domain is a global stretched tripolar ORCA025 grid with an average resolution of $0.25\times 0.25^{\circ}$ which is refined threefold in the Equatorial band. The vertical is resolved with a $z$ coordinate with variable resolution. The upper level is $1\,\mathrm{m}$ thick. This setup is commonly referred to as ORCA025_Z75. The ocean model reads the 10 m wind vectors, temperature, humidity, heat and water fluxes from ERA5 every hour on the native TL639 model resolution, which translates into approximately $31~\mathrm{km}$ horizontal resolution. The ocean model also reads forcing from the ECWAM wave model (a component of ERA5) as described by breivik2015.
3.2 Numerical experiments
We have designed three experiments to isolate the impact of the implementation of the IB effect within a wind-forced system (see Table 1). All experiments cover the period 2015–2022 (inclusively, initialized from a warm start), and none of the runs includes astronomical tides.
Firstly, we perform an ocean-only integration with a constant MSLP (experiment id i5k6) as the control run. We then compute the inverse barometer effect from the ERA5 mean sea level pressure (MSLP) and output the total sea surface elevation (including the inverse barometer effect) as well as the prognostic elevation (only frictional and baroclinic effects). The sea surface elevation from these two experiments are hereafter referred to as noIB for the run using static IB, and diaIB for the run where we compute the inverse barometer effect diagnostically from the local MSLP. The control run is then compared against an otherwise identical ocean-only run where the ocean model is allowed to respond to a varying MSLP taken from ERA5 (experiment id i5k8). This run is referred to as the dynamical experiment (dynIB).
Experiment IB implementation Key physical components
noIB None (i.e. constant MSLP) Wind stress, Ekman transport, baroclinic response
diaIB Diagnostic (post-process) noIB components + diagnostic IB response
dynIB Dynamical (at run-time) Joint pressure-wind dynamics; dynamical IB response
Table 1: Key components of the three experiments conducted to evaluate the two methods of including the IB effect.
In Section 6, we use the IFS ensemble prediction system (ENS, e.g. ens) to showcase how inclusion of the IB effect can impact the probabilistic forecast of storm surges. ENS is a coupled global forecasting system, which uses NEMO for its ocean model component, thus enabling us to explore how the impacts seen from the standalone NEMO studies translate into a coupled forecasting scenario.
Both the uncoupled and coupled (ENS) forecasts use NEMO version 4.0, which is a newer version than the 3.4 version of NEMO which is currently in operational use at ECMWF. The coupled model runs were based on a development version of IFS CY49R2 to incorporate the changes needed to run the coupled setup with NEMO V4.0.
@@ 4.2 Model evaluation and validation
The control experiment with diagnostic inclusion of the IB effect (diaIB) and the alternative (dynIB) is compared with the GESLA3 observations. As a baseline, we also compare against the control experiment without inclusion of the IB effect (noIB). We focus the evaluation on the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE), together with scaled versions of the two ($\mathrm{MAE}/\sigma$ and $\mathrm{RMSE}/\sigma$, where $\sigma$ is the standard deviation of the observed residual water level for each station). The values of MAE and RMSE for each record is scaled by the standard deviation of the observed residual water level for that record to provide a measure of the relative error, since the magnitude of the absolute error tends to be correlated with the magnitude of the residual, to allow us to compare the model performance across different records with different typical amplitudes. This is in line with the approach by kristensen2024, and the formulas can be found in Appendix B.
The MAE and RMSE [m] for the water level of noIB, diaIB and dynIB are listed in Table 2. We focus on the median of the error metrics across the records, but also include the mean values. The median is chosen as it is a robust estimate of central tendency and thus suitable for analysis given the large outliers found in some of the observation records, as explained in Section 2. We include the MAE and RMSE for the entire hindcast period, denoted by the subscript all, and also values where we mask the data when the observed absolute values of the residual is less than one and two standard deviations from the mean, denoted by the subscripts $\pm 1\mathrm{std}$ and $\pm 2\mathrm{std}$. This is done to emphasize the cases and time periods when the contribution from the atmospheric effects to the total water level is significant to focus on the model’s ability to capture anomalously low and high water level events (the latter being what we usually think of as storm surges). Similarly, the relative (scaled) values $\mathrm{MAE}/\sigma$ and $\mathrm{RMSE}/\sigma$, are presented in Table 3.
@@
However, Figure 3 shows that the statistical distribution of the RMSE has a large variation, with the majority of the errors in the range $4-30\,\mathrm{cm}$ and $0.3-1.6]\sigma$, when considering all data. It is evident that all the distributions, except the relative noIB, are skewed to the right. This is as expected based on the mean and median values in Table 2 and 3 where the mean value is generally larger than the median. The width of the distributions increases with stricter masking criteria.
@@
Figure 3: Distribution of the Root Mean Square Error (RMSE) for all records for the three different representations of the atmospheric effects on the water level. The top panels (a-c) present the RMSE in meters, with a bin size of 2 cm. The bottom panels (d-f) show the RMSE scaled by the standard deviation ($\sigma$) of each observation record, using a bin size of $0.1\sigma$. Panels (a) and (d) present the RMSE distribution for the complete dataset. Panels (b) and (e) illustrate the RMSE for observed residual water level absolute values more than one standard deviation ($|\mathrm{obs}|>\sigma$). Panels (c) and (f) show the RMSE for observed residual water level absolute values exceeding two standard deviations ($|\mathrm{obs}|>2\sigma$). The $y$-axis indicates the number of occurrences (records).
Additionally, the map in Figure 4 clearly shows the regional differences in performance between the two experiments. The difference in relative error, displayed in Figure 4, highlights the European NWS and the Baltic Sea as two regions (see also Figure 5) where the differences between the two experiments are large with diaIB outperforming dynIB on the European NWS, and vice versa in the Baltic Sea. Most other areas exhibit relatively small differences, with a few scattered exceptions mostly in favor of the diaIB. It is unclear why the dynIB experiment seems to outperform the diagnostic experiment in the semi-enclosed water body of the Baltic Sea. However, as we see a similar, albeit weaker, tendency in the Mediterranean Sea and the Black Sea, we speculate that the narrow connections to the open ocean limits the efficacy of the inverse barometer effect. These water bodies also have very small tidal amplitudes for the same reason, see e.g., medvedev2016. In other words, although the atmospheric pressure variations set up a pressure gradient, the narrows connecting the Baltic Sea (weisse2021), the Mediterranean Sea and the Black Sea (respectively the Danish Straits, the Strait of Gibraltar and the Bosphorus) are not wide enough for water to adjust on the synoptic time scales of hours to at most a few days. gomis2008 show that for the Mediterranean, the Strait of Gibraltar acts as a choking point and bottleneck where the Atlantic and Mediterranean have independent modes of variability due to the restriction of flow. When the atmospheric pressure averaged over the entire confined sea area changes, this will lead to a change in the average water level in the diaIB, whereas in reality and in the dynIB, these changes in water level are mostly local, leaving the water level over the confined sea area mostly unchanged during the passage of extratropical systems. This effect becomes more pronounced when the dimensions of the confined water body coincides with the spatial extent of the weather systems.
@@
While our global NEMO configuration employs a structured ORCA025 grid ($\sim 31~\mathrm{km}$ resolution), it provides a relatively uniform coastal resolution similar to the 20 km coastal refinement used in the unstructured MOG2D-G mesh of carrere2003. Consistent with their findings at high latitudes, our results show that including the IB effect (whether diagnostic or dynamic) is crucial for reducing errors in energetic regions. However, a key distinction in our study is that wind stress is included in all experiments (noIB, diaIB and dynIB), allowing us to isolate the specific impact of the pressure implementation method.